User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!


When the hydraulics at Ho failed the following error was in the log monitor but was not setting off an alarm so Lucas wasn't woken up for it. It is not listed in the alarm_actions file

WARNING: error fl -1 previous source in this schedule not reached before new source was commanded

Hobart 26m

disk USN-0085, total data pre-recorded 1115.510 GBs

  • Experiment started late, GPIB device continually timed out whilst many errors ensued with control and equip files (Lucas).
  • Turns out experiment was being incorrectly drudged due to unique setup (mk5/dbbc). Jamie manually re-drudged and setupxx worked correctly. First good scan 295-2040 (Lucas).
  • Still getting delay difference changing due to GPIB device timeout.
    • Update, as of 295-2052 after repeated “counter” commands it has fixed itself. Monitoring. (Lucas).
  • I think there is a hydraulics fault. The Live page keeps saying to and keep getting command 'previous source in sched not reached before new source was commanded' (Lucas).
    • Hydraulics fault was caused by circuit break as the motor wasn't covered and water got in. Collaborative effort of Eric and Brett solved the problem. (Lucas).
  • GPIB device continues to time-out, however delay difference is still stable and < 0.3. Counters are being monitored on rack camera to make sure that severe drifts don't occur whilst Eric investigates. (Lucas).
  • 23:50 UT - schedule resumed, lost about 500 GB. First scan 295-2351. GPIB still a problem though, reporting massive delay differences every scan. Clkoff values looks reasonable. (Ellen)
  • 01:04 UT - Eric has reset the GPIB device and it is still a problem. (Ellen)
  • 01:11 UT - Tsys values all overflowing again, restarted CAL again and everything is fine. The CAL has dropped out like this halfway through an experiment before. (Ellen)
  • 01:22 UT - Jamie has turned the GPIB off so that we are no longer measuring a delay difference because it's wrong anyway. Keeping an eye on dbbc=pps_delay instead. (Ellen)
  • 02:17 UT - CAL needed resetting again. And several other times. (Ellen)
  • 07:15 UT - CAL reset again. (Ross)

Hobart 12m

disk HOB+0098, total data pre-recorded 10024.782GBs

  • Experiment Started OK (Lucas).
  • Schedule halted whilst 26m was down. All scans missed between 295-1911 and 295-2016 - ~125 GB missing (Lucas).
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/hob006.txt · Last modified: 2016/10/22 07:16 by Jim Lovell